# An Efficient Dual-Line Decoder Network with Multi-Scale Convolutional Attention for Multi-organ Segmentation

### v1

@@ Abstract
Proper segmentation of organs-at-risk is important for radiation therapy, surgical planning, and diagnostic decision-making in medical image analysis. While deep learning-based segmentation architectures have made significant progress, they often fail to balance segmentation accuracy with computational efficiency. Most of the current state-of-the-art (SOTA) methods either prioritize performance at the cost of high computational complexity or compromise accuracy for efficiency. This paper addresses this gap by introducing an efficient dual-line decoder segmentation network (EDLDNet). The proposed method features a noisy decoder, which learns to incorporate structured perturbation at training time for better model robustness, yet at inference time only the noise-free decoder is executed, leading to lower computational cost. Multi-Scale convolutional Attention Modules (MSCAMs), Attention Gates (AGs), and Up-Convolution Blocks (UCBs) are further utilized to optimize feature representation and boost segmentation performance. By leveraging multi-scale segmentation masks from both decoders, we also utilize a mutation-based loss function to enhance the model’s generalization. Our approach outperforms SOTA segmentation architectures on four publicly available medical imaging datasets (Synapse, ACDC, SegThor, and LCTSC). EDLDNet achieves SOTA performance with an 84.00% Dice score on the Synapse dataset, surpassing baseline model like UNet by 13.89% in Dice score while significantly reducing Multiply-Accumulate Operations (MACs) by 89.7%. Compared to recent approaches like EMCAD, our EDLDNet not only achieves higher Dice score but also maintains comparable computational efficiency. The outstanding performance across diverse datasets establishes EDLDNet’s strong generalization, computational efficiency and robustness. The source code, pre-processed data and pre-trained weights will be available at https://github.com/riadhassan/EDLDNet.
@@
[…]
[C]© 2025. This manuscript version is made available under the CC-BY-NC-ND 4.0 license.
@@
[label4]organization=College of Interdisciplinary Arts and Sciences, Arizona State University, city=Tempe, state=Arizona, country=USA
@@ 1 Introduction
Figure 1: Dice score vs MACs count for different segmentation methods over synapse dataset. The figure presents our method offers the highest segmentation dice score with minimum MACs count.
@@
To overcome these limitations, we propose a segmentation network featuring a dual-line decoder with an efficient multi-scale convolutional attention architecture. Our method introduces a unique training strategy: one decoder operates with imposed noise termed as noisy decoder while the other remains noise-free. During training, both decoders generate segmentation maps in multiple stages that are mutated to create a comprehensive set of combinatorial masks which are used for regularizing the network to increase robustness. In testing or deployment scenarios, only the noise-free decoder is utilized which provides accurate and reliable segmentation with reduced computational overhead. Our architecture achieves superior segmentation performance by demonstrating 13.89% improvement over baseline methods, while 89.7% more computationally efficient (see Figure 1).
@@
    Implement a mutative implicit logit ensemble technique utilizing multi-level logits derived from both noisy and noise free decoders during training. The regularization using mutative logits which enhance the robustness and noise adaptability in segmentation tasks.
@@
    Conduct extensive experiments on four publicly available multi-organ segmentation datasets which show our proposed method outperform state-of-the-art (SOTA) methods. It improves 13.89% in terms of segmentation accuracy (Dice) and 89.7% computation reduction with compared to the baseline. Compared to recent methods, proposed method delivers a higher Dice score while maintaining comparable computational efficiency.
The remainder of the paper is structured as follows: Section 2 highlights on CNN-based and Transformer-based related work; Then, Section 3 presents our Methodology; Experimental details, results of experiments on four publicly available datasets, and ablation studies are discussed in Section 4. Finally, the overall conclusion is drawn in Section 5.
@@ 3 Methodology
In this proposed segmentation network, we utilize a vision transformer based encoder from PVTv2 and two parallel decoder to produce multiple segmentation masks from the network which are then used for mutation. The overall architecture of the proposed method is presented in Figure 2. The following sections outline the encoder and decoder architecture of the segmentation network, the mutation technique employed for optimization during training, and the method for generating the final segmentation mask during inference.
@@ 3.1 Encoder
The PVTv2 architecture [29] enhances traditional transformer patch embedding modules by incorporating consistent spatial information capture with convolutional operations. Utilizing PVTv2-b2 as encoder, we have designed the encoder of our architecture. In our approach, we leverage PVTv2 to extract features labeled as $E_{1}$, $E_{2}$, $E_{3}$ and $E_{4}$ from four stage of the encoder. These features are then fed into decoders, with $E_{4}$ integrated into the upsampling path and $E_{1}$, $E_{2}$ and $E_{3}$ utilized within the skip connections.
@@ 3.2 Decoder
We develop a dual-line decoder with multi-scale convolutional attention to process the extracted features in vision transformer based encoder. Though the dual parallel decoding line are used only in training phase, only one decoder is used in testing or real life applications. The decoders contains Multi-scale Convolutional Attention Modules (MSCAMs) to enhance feature maps, Up-Convolution Blocks (UCBs) for up-sampling and further feature enhancement, and Attention Gates (AGs) to refine the feature maps by integrating skip connections through a gated attention mechanism which are described in Section 3.2.1, and 3.2.2, and 3.2.3 respectively. Both decoders $(j=1,2)$ are identical except a perturbation layer $\mathcal{N}$ is imposed between the encoder and second decoder ($j=2$) presented in Equation 2. Inspired from [9], the perturbation layer consist of uniformly distributed random noise with distribution U(0.3,0.3)U(-0.3,0.3) which is presented in Equation 1.
@@
As the upper stage of decoder carry significant information of segmentation, we consider last two stage decoded features from both decoders (i.e. $D_{i}^{j}$ where, $i=1,2$ and $j=1,2$) followed by $1\times 1$ convolution and bilinear up scaling ($4x$ and $8x$ for $i=1$ and $i=2$ respectively) to match the size of segmentation ground truth mask represented in Equation 5.
@@ 3.2.1 Multi-scale Convolutional Attention Module
To enhance feature refinement, we employ the Multi-scale Convolutional Attention Module (MSCAM) into our network like [21]. It comprised with channel attention, spatial attention, and multi-scale depth-wise convolutions (MSDC) blocks to refine feature maps effectively while maintaining computational efficiency. The MSCAM can be expressed as Equation 7.
@@
MSCB: Multi-Scale Convolution Block is utilized to improve feature representation in the cascaded expanding path. Inspired by the inverted residual block (IRB) architecture introduced in MobileNetV2 [32], the MSCB incorporates a unique approach by employing depth-wise convolutions at varying scales. Additionally, it leverages channel shuffle techniques [33] to facilitate channel mixing across different groups, thereby enhancing the expressiveness and efficiency of the feature extraction process. First the input features are passed through sequentially ($1\times 1$) point wise convolution, batch normalization which is presented in Equation 11 as $S_{mscb}$. After activation with $ReLU$ activation layer, the output features is pass to the three parallel multi scale convolution with karnel size $1\times 1$, $3\times 3$, and $5\times 5$ followed by batch normalization [34] and modified relu $ReLU^{\prime}$ [35] as described in Equation 12. Then the features from three parallel lines are element wise added which is presented in equation 13. Then the output from previous layer is sent to the channel shuffle ($\text{Ch}_{\text{sh}}$) technique [33]. Finally the the features are passed through the sequential layer $S_{mscb}$. The MSCB can be represented by represented by equation 14
@@ 3.2.2 Up Convolution Block
The number of channel reduces in every stage and reversely, during decoding the number of channel increases. The input feature $x_{ch}$ is passed through up convolution for increasing the number of channel where the sequence of channel number $ch\epsilon\{512,320,128,64\}$. The channel number follows the the sequence of Then sequential $3\times 3$ depth wise convolution, batch normalization and $ReLU$ operations are performed to enhance the features. Finally, ($1\times 1$) convolution is performed to reduce the channel to fit this output to the next stage input. The process is presented in Equation 15.
@@ 3.2.3 Attention Gate
Attention gate is implemented in every stage of the decoder to aggregate feature maps where relevant features are activated and suppress the irrelevant features. This attention gate takes two input where one from up-convolution block of the previous stage of decoder and another from corresponding stage skip connection of encoder which are termed as $x_{u}$ and $x_{s}$ respectively. Unlike ($1\times 1$) convolution in Attention-UNet [7], $3\times 3$ grouped convolution [21] is employed for both input $Conv_{3\times 3}(x_{u})$ and $Conv_{3\times 3}(x_{s})$ followed by batch normalization operation is performed which is presented in equation 16. Then element wise addition operation is perform to combine the both features. After that the features are passed through sequential $ReLU$, $1\times 1$ convolution, batch normalization and $Sigmoid$ activation layers which are presented in Equation 17. Output from the sequential operation is element wise multiplied with $x_{u}$ to get the final output from attention module $x_{atten}$. The entire process is expressed as following equation 18.
@@
Where, $\oplus$ represents element wise addition and $\otimes$ represents element wise multipliction.
@@ 3.3 Mutation
Loss aggregation: We employ a combinatorial strategy for loss combination known as mutation [21]. Unlike EMCAD [21], outputs form both decoders (noise-free ($j=1$) and noisy ($j=2$)) are mutated. The mutated features combines both noise-free and noisy properties, thus improving robustness. This method calculates the loss for every possible combination of predictions produced by the four segmentation masks $I$ (see Equation 6), resulting in a total of $2^{4}-1=15$ unique predictions presented in Equation 19. The losses from these combinations are then summed together. As presented in equation 20, we employ loss $\mathcal{L}$ which comprised of Dice loss $\mathcal{L}_{dice}$ and CrossEntropy loss $\mathcal{L}_{ce}$ where weight $\alpha=0.7$ and $\beta=0.3$. Our training process minimizes this cumulative 15 possible combinatorial loss $\mathcal{L}_{total}$ presented in Equation 21.
$\mathcal{P}(I)=\{A\mid A\subseteq I,A\neq\emptyset,\text{Sum}(A)=\sum_{M\in A}M\}$ (19)
@@
$\mathcal{L}_{total}=\sum_{i=1}^{15}\mathcal{L}(softmax(\mathcal{P}_{i}),target)$ (21)
@@ 3.4 Final Segmentation Mask Generation:
For inference or real-time applications, we utilize only the prediction mask $I_{1}^{1}$ as the final output where only the encoder and the noise-free decoder ($j=1$) is necessary. Although the noisy decoder ($j=2$) is employed during training to guide the noise-free decoder towards improved robustness and noise adaptability, it is unnecessary during testing, deployment or real-time applications. The final segmentation results are generated by applying the Softmax function for multi-class segmentation, as described in equation 22.
@@ 4.1 Dataset
To evaluate the effectiveness of our proposed method, we use four publicly available multi-organ segmentation datasets including Synapse [36]  [^1] https://doi.org/10.7303/syn3193805, ACDC  [^2] https://www.creatis.insa-lyon.fr/Challenge/acdc/index.html [37], SegThor  [^3] https://competitions.codalab.org/competitions/21145 [38], and LCTSC  [^4] https://doi.org/10.7937/K9/TCIA.2017.3R3FVZ08 [39].
@@ 4.2 Experimental Parameters
To conduct our experiment, we used 300 epoch for synapse, and ACDC and 200 epoch for SegThor, and LCTSC in train and the batch size was 6. For optimization, we use AdamW optimizer with learning rate 0.0001 and weight decay 0.0001. The input size is $224\times 224$ for Synapse, and ACDC dataset and $256\times 256$ for SegThor, and LCTSC dataset. To evaluate the performance of different segmentation methods, we utilize popular metrics Dice score, mean Intersection over Union (mIoU), Asymmetric Surface Distance (ASD) [40]. For evaluating the required computation of a segmentation method, we use Multiply-Accumulate Operations (MACs) count [41].
@@ 4.4 Results on Synapse Dataset
Among the earlier models, UNet serves as a baseline, achieving a Dice score of 70.11%. Its improved version, Attention UNet, incorporates attention mechanisms to focus on relevant spatial regions, slightly boosting performance to 71.70%. However, these models struggle with segmenting smaller and more complex organs like the pancreas and gallbladder. The introduction of hybrid architectures such as R50+UNet and R50+Attention UNet further enhances performance, pushing Dice scores to 74.68% and 75.57%, respectively. Despite these improvements, the increased computational cost is a major drawback—R50+Attention UNet requires 91.72G MACs, making it impractical for real-time applications. Transformer-based models have significantly transformed medical image segmentation by leveraging self-attention mechanisms to capture long-range dependencies. TransUNet, one of the first Transformer-based models in this domain, achieves a notable jump in performance with a Dice score of 77.61%, while being computationally more efficient than its ResNet-based UNet counterparts (32.46G MACs). More recent Transformer-based architectures, such as SSFormer, PolypPVT, and MT-UNet, continue this trend of improvement. Notably, MT-UNet achieves a Dice score of 78.96%, while PolypPVT delivers a similar 78.08% with lower computational demands. UDBANet utilizes uncertainty information within the network and it achieves 79.99% dice score with 24.59G MACs. Among these, PVT-CASCADE stands out, reaching an impressive Dice score of 81.06% while keeping MACs at just 8.12G, highlighting the power of PVT in achieving both high accuracy and efficiency. At the top end of the performance spectrum, GCASCADE, EMCAD, and our proposed model achieve SOTA segmentation results while maintaining efficiency. GCASCADE delivers a Dice score of 83.28%, while EMCAD refines it further to 83.63%, with an improved mIoU of 74.65%. Our proposed model, EDLDNet, achieves SOTA performance with an 84.00% Dice score and 75.03% mIoU, surpassing all previous methods. Despite this significant performance improvement, the model remains highly computationally efficient, requiring only 5.6G MACs, making it well-suited for real-world applications. EDLDNet outperforms baseline methods like UNet by 20% in accuracy while reducing computational cost by 89% in terms of MACs Operations. Compared to recent approaches such as EMCAD, EDLDNet demonstrates marginal yet consistent improvements in segmentation accuracy while maintaining comparable efficiency, underscoring its ability to balance high performance with low computational complexity.
@@
At the organ level performance, our model achieves the highest Dice scores for the aorta (89.12%), gallbladder (74.15%), liver (95.56%), and stomach (83.99%), demonstrating its robustness in segmenting both large and small organs. While EMCAD performs slightly better on the pancreas (68.51%) and spleen (92.17%), our model remains highly competitive, achieving 67.89% and 91.34%, respectively. Kidney segmentation remains particularly challenging, with EMCAD achieving the best Dice score for the left kidney (88.08%), closely followed by our model at 87.27%. For the right kidney, TransCASCADE edges out our model (84.56% vs. 82.63%), indicating areas where further refinements could improve performance. Our model strikes an ideal balance, achieving SOTA accuracy while maintaining an extremely low computational requirement as presented in Figure 1 which makes it a viable solution for practical applications where both precision and efficiency are critical.

### v2

@@ Abstract
Proper segmentation of organs-at-risk is important for radiation therapy, surgical planning, and diagnostic decision-making in medical image analysis. While deep learning-based segmentation architectures have made significant progress, they often fail to balance segmentation accuracy with computational efficiency. Most of the current state-of-the-art (SOTA) methods either prioritize performance at the cost of high computational complexity or compromise accuracy for efficiency. This paper addresses this gap by introducing an efficient dual-line decoder segmentation network (EDLDNet). In addition to noise-free decoder, the proposed method features a noisy decoder, which learns to incorporate structured perturbation at training time for better model robustness, yet at inference time only the noise-free decoder is executed, leading to lower computational cost. Multi-Scale Convolutional Attention Modules (MSCAMs), Attention Gates (AGs), and Up-Convolution Blocks (UCBs) are further utilized to optimize feature representation and boost segmentation performance. By leveraging multi-scale segmentation masks from both decoders, a mutation-based loss function is also utilized to enhance the model’s generalization. The proposed method outperforms SOTA segmentation architectures on four publicly available medical imaging datasets (Synapse, ACDC, SegThor, and LCTSC). EDLDNet achieves SOTA performance with an 84.00% Dice score on the Synapse dataset, surpassing baseline model like UNet by 13.89% in Dice score while significantly reducing Multiply-Accumulate Operations (MACs) by 89.7%. Compared to recent approaches like EMCAD, the proposed EDLDNet not only achieves higher Dice score but also maintains comparable computational efficiency. The outstanding performance across diverse datasets establishes EDLDNet’s outstanding generalization, computational and robustness efficiency. The source code, pre-processed data, and pretrained weights are available at https://github.com/riadhassan/EDLDNet.
@@
[label4]organization=New College of Interdisciplinary Arts and Sciences, Arizona State University, city=Tempe, state=Arizona, country=USA
@@ 1 Introduction
Figure 1: Dice score vs MACs count for different segmentation methods over synapse dataset. The figure presents that the proposed method offers the highest segmentation dice score with minimum MACs count.
@@
To overcome these limitations, we propose a segmentation network featuring a dual-line decoder with an efficient multi-scale convolutional attention architecture. This method introduces a unique training strategy: one decoder operates with imposed noise termed as noisy decoder while the other remains noise-free. During training, both decoders generate segmentation maps in multiple stages that are mutated to create a comprehensive set of combinatorial masks which are used for regularizing the network to increase robustness. In testing or deployment scenarios, only the noise-free decoder is utilized which provides accurate and reliable segmentation with reduced computational overhead. This architecture achieves superior segmentation performance by demonstrating 13.89% improvement over baseline methods, while 89.7% more computationally efficient (see Figure 1).
@@
    Implement a mutative implicit logit ensemble technique utilizing multi-level logits derived from both noisy and noise-free decoders during training. The regularization with mutative logits enhances robustness and noise adaptability in segmentation tasks.
@@
    Conduct extensive experiments on four publicly available multi-organ segmentation datasets which show the proposed method outperform state-of-the-art (SOTA) methods. It improves 13.89% in terms of segmentation accuracy (Dice) and 89.7% computation reduction with compared to the baseline. Compared to recent methods, proposed method delivers a higher Dice score while maintaining comparable computational efficiency.
The remainder of the paper is structured as follows: Section 2 highlights on CNN-based and Transformer-based related work; then, Section 3 presents Methodology. Experimental details, results of experiments on four publicly available datasets, and ablation studies are discussed in Section 4. Finally, the overall conclusion is drawn in Section 5.
@@ 3 Methodology
In this proposed segmentation network, we employs a vision transformer-based encoder from PVTv2, coupled with dual parallel decoders—a noisy decoder and a noise-free decoder. Each decoder integrates Multi-scale Convolutional Attention Modules (MSCAMs), Up-Convolution Blocks (UCBs), and Attention Gates (AGs) to enhance feature representation and segmentation accuracy. Unlike the noise-free decoder, the noisy decoder incorporates an additional noise layer in its architecture. During training, both decoders generate multiple segmentation logits for mutation-based regularization, while only the noise-free decoder is active during inference to ensure computational efficiency. The complete architecture of the proposed method is illustrated in Figure 2. The subsequent sections detail the encoder and decoder designs, the mutation strategy for training optimization, and the process for generating the final segmentation mask during inference.
@@ 3.1 Encoder
The proposed architecture employs the PVTv2-B2 transformer as the encoder, chosen for its efficient hierarchical feature extraction capabilities and less computation requirement which are essential for efficient medical image segmentation [29]. The encoder initiates processing through an overlapping patch embedding layer using a $4\times 4$ convolutional kernel with stride 4. The four-stage hierarchical architecture is configured with embedding dimensions of $[64,128,320,512]$, utilizing $[1,2,5,8]$ attention heads and $[3,4,6,3]$ transformer blocks per stage, respectively. Each stage incorporates spatial-reduction attention (SRA) with reduction ratios of $[8,4,2,1]$. The MLP layers employ expansion ratios of $[8,8,4,4]$, with Layer Normalization ($\epsilon=1\times 10^{-6}$) and learnable QKV biases. Stochastic depth regularization (drop path rate = 0.1) is applied while disabling standard dropout (rate = 0.0) for optimal feature learning. This configuration generates multiscale features labeled as $E_{1}$, $E_{2}$, $E_{3}$ and $E_{4}$ from four stage of the encoder. These features are then fed into decoders, with $E_{4}$ integrated into the upsampling path and $E_{1}$, $E_{2}$ and $E_{3}$ utilized within the skip connections.
@@ 3.2 Decoder
A dual-line decoder with multi-scale convolutional attention is developed to process the extracted features in vision transformer based encoder. Though the dual parallel decoding line are used only in training phase, only one decoder is used in testing or real life applications. The decoders contains MSCAMs to enhance feature maps, UCBs for up-sampling and further feature enhancement, and AGs to refine the feature maps by integrating skip connections through a gated attention mechanism which are described in Section 3.2.1, and 3.2.2, and 3.2.3 respectively. Both decoders $(j=1,2)$ are identical except a perturbation layer $\mathcal{N}$ is imposed between the encoder and second decoder ($j=2$) presented in Equation 2. Inspired from [9], the perturbation layer consist of uniformly distributed random noise with distribution U(0.3,0.3)U(-0.3,0.3) which is presented in Equation 1. The architecture of the noise free and noisy decoder are presented in Figure 2 (b) and Figure 2 (c) respectively.
@@
As the upper stage of decoder carry significant information of segmentation, last two stage decoded features are considered from both decoders (i.e. $D_{i}^{j}$ where, $i=1,2$ and $j=1,2$) followed by $1\times 1$ convolution and bilinear up scaling ($4x$ and $8x$ for $i=1$ and $i=2$ respectively) to match the size of segmentation ground truth mask represented in Equation 5.
@@ 3.2.1 Multi-scale Convolutional Attention Module
To enhance feature refinement, the Multi-scale Convolutional Attention Module (MSCAM) is employed into the proposed network like [21]. It comprised with channel attention, spatial attention, and multi-scale depth-wise convolutions (MSDC) blocks to refine feature maps effectively while maintaining computational efficiency. The MSCAM can be expressed as Equation 7. The architecture of the MSCAM is illustrated in Figure 2 (e).
@@
MSCB: Multi-Scale Convolution Block is utilized to improve feature representation in the cascaded expanding path. Inspired by the inverted residual block (IRB) architecture introduced in MobileNetV2 [32], the MSCB incorporates a unique approach by employing depth-wise convolutions at varying scales. Additionally, it leverages channel shuffle techniques [33] to facilitate channel mixing across different groups, thereby enhancing the expressiveness and efficiency of the feature extraction process. First the input features are passed through sequentially ($1\times 1$) point wise convolution, batch normalization which is presented in Equation 11 as $S_{mscb}$. After activation with $ReLU$ activation layer, the output features is pass to the three parallel multi scale convolution with karnel size $1\times 1$, $3\times 3$, and $5\times 5$ followed by batch normalization [34] and modified relu $ReLU^{\prime}$ [35] as described in Equation 12. Then the features from three parallel lines are element wise added which is presented in equation 13. Then the output from previous layer is sent to the channel shuffle ($\text{Ch}_{\text{sh}}$) technique [33]. Finally the the features are passed through the sequential layer $S_{mscb}$. The MSCB can be represented by equation 14.
@@ 3.2.2 Up Convolution Block
The number of channel reduces in every stage and reversely, during decoding the number of channel increases. The input feature $x_{ch}$ is passed through up convolution for increasing the number of channel where the sequence of channel number $ch\epsilon\{512,320,128,64\}$. The channel number follows the the sequence of Then sequential $3\times 3$ depth wise convolution, batch normalization and $ReLU$ operations are performed to enhance the features. Finally, ($1\times 1$) convolution is performed to reduce the channel to fit this output to the next stage input. The process is presented in Equation 15 and the architecture is presented in Figure 2 (f).
@@ 3.2.3 Attention Gate
Attention gate is implemented in every stage of the decoder to aggregate feature maps where relevant features are activated and suppress the irrelevant features. This attention gate takes two input where one from up-convolution block of the previous stage of decoder and another from corresponding stage skip connection of encoder which are termed as $x_{u}$ and $x_{s}$ respectively. Unlike ($1\times 1$) convolution in Attention-UNet [7], $3\times 3$ grouped convolution [21] is employed for both input $Conv_{3\times 3}(x_{u})$ and $Conv_{3\times 3}(x_{s})$ followed by batch normalization operation is performed which is presented in equation 16. Then element wise addition operation is perform to combine the both features. After that the features are passed through sequential $ReLU$, $1\times 1$ convolution, batch normalization and $Sigmoid$ activation layers which are presented in Equation 17. Output from the sequential operation is element wise multiplied with $x_{u}$ to get the final output from attention module $x_{atten}$. The entire process is expressed as following equation 18. The architecture of the attention gate is presented in Figure 2 (d).
@@
Where, $\oplus$ represents element wise addition and $\otimes$ represents element wise multiplication.
@@ 3.3 Mutation
Loss aggregation: We employ a combinatorial strategy for loss combination known as mutation [21]. Unlike EMCAD [21], outputs from both decoders (noise-free ($j=1$) and noisy ($j=2$)) are mutated. The mutated features combine both noise-free and noisy properties, thus improving robustness. This method calculates the loss for every possible non-empty subset of predictions produced by the four segmentation logits $I$ (see Equation 6), resulting in a total of $|\mathcal{P}(I)|=2^{4}-1=15$ unique predictions as defined in Equation 19. The losses from these combinations are then summed together. As presented in Equation 20, the loss function $\mathcal{L}(\cdot)$ combines Dice loss $\mathcal{L}_{dice}(\cdot)$ and CrossEntropy loss $\mathcal{L}_{ce}(\cdot)$ with weights $\alpha=0.7$ and $\beta=0.3$. This training process minimizes the cumulative combinatorial loss $\mathcal{L}_{\text{total}}$ defined in Equation 21.
$\mathcal{P}(I)=\{A\mid A\subseteq I,A\neq\emptyset\}$ (19)
@@
$\mathcal{L}_{\text{total}}=\sum_{i=1}^{15}\mathcal{L}\left(\text{softmax}\left(\sum_{x\in P_{i}}x\right),\text{target}\right)$ (21)
@@ 3.4 Final Segmentation Mask Generation:
For inference or real-time applications, only the prediction mask $I_{1}^{1}$ is utilized as the final output where only the encoder and the noise-free decoder ($j=1$) is necessary. Although the noisy decoder ($j=2$) is employed during training to guide the noise-free decoder towards improved robustness and noise adaptability, it is unnecessary during testing, deployment or real-time applications. The final segmentation results are generated by applying the Softmax function for multi-class segmentation, as described in equation 22.
@@ 4.1 Dataset
To evaluate the effectiveness of the proposed method, we use four publicly available multi-organ segmentation datasets including Synapse [36]  [^1] https://doi.org/10.7303/syn3193805, ACDC  [^2] https://www.creatis.insa-lyon.fr/Challenge/acdc/index.html [37], SegThor  [^3] https://competitions.codalab.org/competitions/21145 [38], and LCTSC  [^4] https://doi.org/10.7937/K9/TCIA.2017.3R3FVZ08 [39].
@@ 4.2 Experimental Parameters
To conduct experiments, we used 300 epoch for synapse, and ACDC and 200 epoch for SegThor, and LCTSC in train and the batch size was 6. For optimization, AdamW optimizer are used with learning rate 0.0001 and weight decay 0.0001. The input size is $224\times 224$ for Synapse, and ACDC dataset and $256\times 256$ for SegThor, and LCTSC dataset. To evaluate the performance of different segmentation methods, we utilize popular metrics Dice score, mean Intersection over Union (mIoU), Asymmetric Surface Distance (ASD) [40]. For evaluating the required computation of a segmentation method, we use Multiply-Accumulate Operations (MACs) count [41].
@@ 4.4 Results on Synapse Dataset
Among the earlier models, UNet serves as a baseline, achieving a Dice score of 70.11%. Its improved version, Attention UNet, incorporates attention mechanisms to focus on relevant spatial regions, slightly boosting performance to 71.70%. However, these models struggle with segmenting smaller and more complex organs like the pancreas and gallbladder. The introduction of hybrid architectures such as R50+UNet and R50+Attention UNet further enhances performance, pushing Dice scores to 74.68% and 75.57%, respectively. Despite these improvements, the increased computational cost is a major drawback—R50+Attention UNet requires 91.72G MACs, making it impractical for real-time applications. Transformer-based models have significantly transformed medical image segmentation by leveraging self-attention mechanisms to capture long-range dependencies. TransUNet, one of the first Transformer-based models in this domain, achieves a notable jump in performance with a Dice score of 77.61%, while being computationally more efficient than its ResNet-based UNet counterparts (32.46G MACs). More recent Transformer-based architectures, such as SSFormer, PolypPVT, and MT-UNet, continue this trend of improvement. Notably, MT-UNet achieves a Dice score of 78.96%, while PolypPVT delivers a similar 78.08% with lower computational demands. UDBANet utilizes uncertainty information within the network and it achieves 79.99% dice score with 24.59G MACs. Among these, PVT-CASCADE stands out, reaching an impressive Dice score of 81.06% while keeping MACs at just 8.12G, highlighting the power of PVT in achieving both high accuracy and efficiency. At the top end of the performance spectrum, GCASCADE, EMCAD, and the proposed model achieve SOTA segmentation results while maintaining efficiency. GCASCADE delivers a Dice score of 83.28%, while EMCAD refines it further to 83.63%, with an improved mIoU of 74.65%. The proposed model, EDLDNet, achieves SOTA performance with an 84.00% Dice score and 75.03% mIoU, surpassing all previous methods. Despite this significant performance improvement, the model remains highly computationally efficient, requiring only 5.6G MACs, making it well-suited for real-world applications. EDLDNet outperforms baseline methods like UNet by 20% in accuracy while reducing computational cost by 89% in terms of MACs Operations. Compared to recent approaches such as EMCAD, EDLDNet demonstrates marginal yet consistent improvements in segmentation accuracy while maintaining comparable efficiency, underscoring its ability to balance high performance with low computational complexity.
@@
At the organ level performance, the proposed model achieves the highest Dice scores for the aorta (89.12%), gallbladder (74.15%), liver (95.56%), and stomach (83.99%), demonstrating its robustness in segmenting both large and small organs. While EMCAD performs slightly better on the pancreas (68.51%) and spleen (92.17%), our model remains highly competitive, achieving 67.89% and 91.34%, respectively. Kidney segmentation remains particularly challenging, with EMCAD achieving the best Dice score for the left kidney (88.08%), closely followed by proposed model at 87.27%. For the right kidney, TransCASCADE edges out our model (84.56% vs. 82.63%), indicating areas where further refinements could improve performance. This model strikes an ideal balance, achieving SOTA accuracy while maintaining an extremely low computational requirement as presented in Figure 1 which makes it a viable solution for practical applications where both precision and efficiency are critical.